### v1

# Neural Network-Based Bandit: A Medium Access Control for the IIoT Alarm Scenario

Abstract
Efficient Random Access (RA) is critical for enabling reliable communication in Industrial Internet of Things (IIoT) networks. Herein, we propose a deep reinforcement learning based distributed RA scheme, entitled Neural Network-Based Bandit (NNBB), for the IIoT alarm scenario. In such a scenario, the devices may detect a common critical event, and the goal is to ensure the alarm information is delivered successfully from at least one device. The proposed NNBB scheme is implemented at each device, where it trains itself online and establishes implicit inter-device coordination to achieve the common goal. Devices can transmit simultaneously on multiple orthogonal channels and each possible transmission pattern constitutes a possible action for the NNBB, which uses a deep neural network to determine the action. Our simulation results show that as the number of devices in the network increases, so does the performance gain of the NNBB compared to the Multi-Armed Bandit (MAB) RA benchmark. For instance, NNBB experiences a 7% success rate drop when there are four channels and the number of devices increases from 10 to 60, while MAB faces a 25% drop.
Index Terms:
Alarm scenario, Deep Reinforcement Learning, Industrial Internet of Things, Multi-Armed Bandit, Neural Network-Based Bandit, Random Access.
I INTRODUCTION
The Industrial Internet of Things (IIoT) incorporates intelligence and autonomy into manufacturing through dependable wireless connectivity, production line insights, smart alerts, and predictive maintenance via data acquisition and processing from numerous devices. Notably, the number of IIoT devices is expected to reach around $9.4$ billion by $2030$, representing $37\%$ of the total population of devices [1], and will lead to massive IIoT networks.
Some critical challenges in massive IIoT networks include wireless channel overload, packet loss due to high latency and deadline violation, and energy consumption escalation at the devices, e.g., due to repeated channel access requests. These challenges constitute the so-called massive access problem [2, 3]. In general, channel access solutions addressing this problem in IIoT deployments must adapt to sporadic and correlated traffic [4, 5], require low signaling [6, 5], support decentralized decision-making [7, 5], and be scalable [7, 5]. Orthogonal Multiple Access (MA) schemes cannot be used since they require excessive signaling, e.g., long orthogonal pilot sequences, and do not support decentralized decision-making. Although Aloha-based Random Access (RA) schemes satisfy the first three conditions, they are plagued by limited scalability, making them unfeasible in a massive IIoT network. This, and the boom of tinyML for on-device intelligence [8, 9], has motivated the research on learning-based RA for massive IIoT networks.
I-A STATE-OF-ART ON LEARNING-BASED RA
Several learning-based RA schemes have already been proposed in the literature. As shown in Table I, Reinforcement Learning (RL) [10, 11, 12, 13], Deep RL (DRL) [14, 7, 15], and Deep Q-Network (DQN) [16, 17, 18, 19] are the most popular methods adopted by the existing learning-based RA works.
The authors in [10] proposed a RL-based distributed RA scheme for Delay-Constrained (DC) communications (RLRA-DC). It establishes cooperation among devices to increase the system throughput beyond the maximum system throughput of $1/e$ achieved by the Aloha-based schemes. However, RLRA-DC requires information about the total number of devices present in the network. Although there are approaches to estimate such information, e.g., [10, 20], they are by no means error-free, inevitably leading to performance degradation of RLRA-DC in practice. The authors in [11] proposed a Tiny State-space R-learning random Access (TSRA) scheme to convey DC traffic. TSRA is a distributed RA scheme exploiting information regarding the urgency level of the data packets, which is measured in terms of their time to expire. The most important feature of TSRA is the independence of the size of the device state space from the number of devices in the network and the hard deadline for transmitting a packet. In spite of this, RLRA-DC achieves a higher system throughput than TSRA.
In [12], the authors considered a multi-access point (AP) scenario, where each user runs a two-stage RL based RA algorithm for maximizing the throughput. In the first stage, the user selects the AP at the beginning of the macro-time slot, while in the second stage, the user decides the sub-time slot for transmission. Despite its appealing distributed nature, the proposal has the following limitations: (i) a large number of devices in the network would make the lookup table of the second RL algorithm too large, thus taking a significant time to converge and degrading the solution’s optimality; and (ii) information about the devices associated with an AP is required after every macro-time slot.
The authors in [13] proposed a coordinated RA scheme adapted to sporadic and correlated IIoT traffic. Therein, the time slot selection task at each device is modelled as a Markov game, where devices are the agents, the number of retransmissions is the context, and the transmit time slots in a frame are the actions. Moreover, each device runs an RL-based Linear Reward-Inaction (LRI) algorithm [21] to learn the equilibrium points of the Markov game. Despite its appealing distributed nature, LRI in [13] never converged to a pure Nash equilibrium, leading to the coordinated RA scheme being beneficial only for moderate traffic conditions. Meanwhile, the authors in [14] developed a type of contention-cum-RL-based RA scheme called Inner-State-Driven random Access (ISDA). ISDA uses a Neural Network (NN) with one hidden layer of five neurons to determine the transmission probabilities of the devices in a time slot as a function of their inner states. The main benefits of ISDA include the decentralized modus operandi and the ability to meet heterogeneous performance requirements. However, its scalability is unproven as it was tested for a network with only three devices.
TABLE I: Existing Learning-based RA Research Works
Ref. Learning method Scenario
[13] LRI Uplink transmission of correlated IIoT traffic through a cellular network
[6] MAB Transmission of sensed alarm information to the ExC
[7] DRL Uplink transmission in a massive IoT network
[10, 11] R-learning Uplink transmission of delay-constrained traffic
[12] MAB Uplink transmission in a multicell system
[14] Cross-entropy [22] Network having a diverse quality of service demands from multiple IoT applications
[15] PPO Uplink transmission in heterogeneous IoT network
[16] DQN Coexisting networks, each operating with a distinct medium access protocol
[17] DQN Uplink transmission of sporadic traffic on a single channel in a slotted system
[18] DQN Millimeter-wave network with numerous small/macro cells, consisting of traversing users
[19] DQN Efficient spectrum access in a multichannel network with a simple collision model
The authors in [7] proposed a multi-agent DRL-based RA scheme that employs a centralized training and decentralized execution approach for the uplink transmission in massive IoT setups. The learned RA policy in [7] is scalable, it can be shared among all the devices, and is adaptable to both correlated and regular IoT traffic. However, it has two flaws: (i) high complexity as devices must possess a NN with $256$ neurons in each of its two hidden layers, and (ii) convergence issues. To enhance the overall network throughput of a heterogeneous IoT network, the authors in [15] proposed a distributed Proximal Policy Optimization (PPO)-based MA (PPOMA). IoT nodes adopting PPOMA perform online learning to update their access policies and dynamically coexist with other nodes. A notable advantage of PPOMA is that it operates without demanding information about the MA protocols employed by the other coexisting IoT nodes. However, PPOMA requires three distinct NNs, namely actor, critic, and actor-target NNs, to be stored in the IoT node’s memory. Additionally, [15] acknowledges that while PPOMA enhances the overall network throughput, it does so at the expense of fairness among the coexisting IoT nodes.
As mentioned earlier, DQN has also been adopted by many works. In [16], the authors introduced a DRL MA (DLMA) scheme exploiting DQN and aiming to maximize the sum throughput and the $\alpha$-fairness among the coexisting networks. DLMA allows the devices to learn how to properly use the time-spectral resources with no information about the medium access protocols used by other coexisting networks. However, DLMA demands an extensive NN with six hidden layers and $64$ neurons per each, and still cannot outperform the R-learning-based TSRA in terms of throughput, as stated in [11]. To learn a transmission policy that balances out throughput with fairness among users, the authors in [17] proposed a multi-agent DQN RA scheme. Here, fairness is measured via the age of a packet. The scheme employs a parameter-sharing method from [23] to train just one NN in a centralized manner and extend it to work for all devices. Unfortunately, even though the DQN RA scheme in [17] can tune itself to different Poisson-distributed data arrival rates, its performance degrades as the number of devices increases, as stated in [24].
The authors in [18] proposed a DQN-based algorithm for wandering users to estimate the congestion levels of APs and select them accordingly, thus reducing the delay experienced by the users during contention-based RA. Meanwhile, the total number of successfully delivered packets was maximized in [19] by solving a dynamic spectrum access problem in a multichannel network. Because of the availability of many possible network states and the partial observability of those states due to zero message exchange among the devices, a pre-trained DQN-based RA scheme is proposed. The channel selection decisions are taken online and in a distributed fashion by the devices. However, a serious limitation is that the DQN at each device needs to be retrained offline every week or every time a significant environmental change occurs. It is worth noting that the DQN-based works [16, 18, 17, 19] have one limitation in common, which is that they all require a replica of the former NN model (also called target model) for training the latest NN model. Specifically, the approaches in [16, 18] require a replica to be stored in the device’s memory, while those in [17, 19] require it to be stored at the central unit.
I-B MOTIVATION AND CONTRIBUTION
As discussed above, there are numerous learning-based RA schemes in the recent literature addressing several wireless communication scenarios. However, the IIoT alarm scenario remains much less explored. The alarm scenario consists of IoT devices that may sense the same critical event, such as an abnormal turbine vibration or an oil leak detection, and must promptly inform an External Controller (ExC). Hence, they must adopt a low-overhead protocol that avoids collisions from simultaneous alarm transmissions so that the critical alarm message transmission from at least one device succeeds.
TABLE II: List of Symbols
Symbol Description
$\boldsymbol{\varphi}$, $\boldsymbol{\hat{\varphi}}$ Additive white Gaussian noise vector
$\rho$ Average transmit signal-to-noise ratio
$\Lambda$ Expected probability of successful transmission
$\Omega$ Loss function
$\tau$ Learning rate for MAB
$\lambda$ Mean-scaling multiplier
$\gamma$ Path loss exponent
$\varepsilon$ Probability of random action selection
$\beta_{0}$ Threshold value for ${\|\nabla_{\mathbf{w}}\Omega\|}_{2}$
$\boldsymbol{\chi}$ Vector for storing the clipped gradients
$\mathcal{N}^{\prime}$ Batch of active devices
$\mathcal{N}$ Set of $N$ devices
$\mathbf{w}$ DNN weight vector
$\mathbf{\mathring{A}}$ Matrix containing all possible transmission patterns
$\mathbf{\Psi}$ Matrix containing the probability of each device selecting a transmission pattern
$\mathbf{A}$ Matrix storing the transmission patterns chosen by the active devices
$\boldsymbol{\mathring{a}}$ Transmission pattern
$\boldsymbol{s}$ Aggregated pilot signal received by the BS
$E$ Finite memory buffer
$M$ Number of available orthogonal channels
$H$ Number of hidden layers
$h$ Size of each hidden layer
$B$ Size of mini-batch for training the DNN
$\boldsymbol{a}_{\upsilon}$ Action selected by the active agent $\upsilon$
$\mathbf{c}_{\upsilon}$ Channel coefficients between device $\upsilon$ and the BS
$\boldsymbol{s}_{\upsilon}$ Context received by the active device $\upsilon$
$d_{\upsilon}$ Distance between device $\upsilon$ and the alarm epicentre
$r_{\upsilon}$ Distance between device $\upsilon$ and the BS
$\boldsymbol{\varrho_{\upsilon}}$ Pilot sequence for an active agent $\upsilon$
$\varrho_{i,\upsilon}$ $i^{th}$ pilot symbol of the active agent $\upsilon$
$f(d_{\upsilon})$ Activation probability function of device $\upsilon$
$Q(\boldsymbol{\mathring{a}}_{i})$ Action value of $\boldsymbol{\mathring{a}}_{i}$ in case of MAB
$\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})$ Probability of device $\upsilon$ choosing $\boldsymbol{\mathring{a}}_{i}$
$\xi(\mathcal{N}^{\prime},\mathbf{A})$ Indicator for a successful///unsuccessful transmission
$\hat{q}(\boldsymbol{s},\boldsymbol{\mathring{a}}_{i},\mathbf{w})$ Parameterized action-value of $\boldsymbol{\mathring{a}}_{i}$ given $\boldsymbol{s}$ and $\mathbf{w}$
$r(\boldsymbol{a}_{\upsilon})$ Reward received by active agent $\upsilon$ for selecting $\boldsymbol{a}_{\upsilon}$
$\nabla_{\mathbf{w}}\Omega$ Gradient of $\Omega$ with respect to $\mathbf{w}$
In this work, we focus on the alarm scenario in an IIoT network. This problem has been studied earlier in [6], where the authors proposed a Multi-Armed Bandit (MAB) based RA scheme to allow the active devices, which detect the alarm, to indirectly coordinate their alarm transmissions to the ExC. The main limitations of the approach in [6] are: (i) no more than three devices can be active concurrently; (ii) training occurs at the active devices but only one at a time, which is time-consuming and imposes tight synchronization requirements on the network; and (iii) MAB does not perform well with a large action space. Meanwhile, existing Industrial Wireless Sensor Network (IWSN) protocols such as WirelessHART [25], ISA100.11100.11100.11100.11a [25], WIA-PA [25], employ Time Division MA (TDMA). TDMA is favored because it allows the prediction of communication latency for time-triggered packets in advance. However, centralized approaches like TDMA struggle to deliver event-triggered packets, such as IIoT alarm message packets, in a timely manner since the generation of the event-triggered packets is unpredictable [26]. The limitations of [6] and existing IWSN protocols motivate our work, which presents a novel, efficient, online learning-based RA for the alarm scenario. Our contributions are as follows:
  • We devise a novel procedure to acquire a useful context to assist the learning-based RA. The procedure starts with the active devices transmitting their respective pilot signals to a Base Station (BS). These signals are received as an aggregated signal at the BS, which then broadcasts it to the active devices. At each device, this constitutes the context for the proposed RL framework.
  • We propose a distributed learning-based RA scheme called Neural Network-Based Bandit (NNBB) for the IIoT alarm scenario. NNBB uses a simple Deep NN (DNN) to process the context received by a device and decide the transmission channel(s) for the alarm signals such that the ExC successfully receives the signal on at least one channel. Unlike [6], there is no restriction on the number of simultaneously active devices.
  • We devise an online training procedure for the NNBB in which the DNN weights are updated at every alarm event using Root Mean Square Propagation (RMSProp). Unlike DQN, NNBB does not require a target model for training.
  • We compare NNBB with two benchmark schemes, namely MAB-based RA and Random Selection (RS), under various network configurations through simulations. Our results indicate that NNBB outperforms them even when the network experiences an increase in the available channels/devices or activation probability.
Figure 1: IIoT alarm scenario. An alarm event at a certain location triggers some nearby devices, which then become active and must convey their sensed alarm information to the ExC. Here, the BS broadcasts the context, which is the aggregated pilot signal, to the active devices.
I-C ORGANIZATION
The paper is structured as follows. Section  II introduces the system model and problem formulation. Section  III describes the proposed NNBB scheme and the Medium Access Control (MAC) procedure followed by an IIoT device. We analyze the performance of NNBB through simulations in Section  IV. Finally, we conclude the paper in Section  V and also highlight some open research areas.
Notation: $\operatorname{argmax}(\cdot)$ and $\max(\cdot)$ are the argument of the maximum function and the maximum function, respectively. $\mathcal{P}(\cdot)$ is the power set and ||$\cdot$|| ⋅ | represents the cardinality of a set. Superscript $[\cdot]^{T}$ is the transpose operation and ${\|\cdot\|}_{2}$ is the Euclidean norm of a vector. Column vectors/matrices are denoted by boldface lowercase/uppercase letters, and $\mathbf{I}_{M}$ is the $M\times M$ identity matrix. $\mathbb{C}^{M\times 1}$ is a set of complex vectors of dimension $M\times 1$. A circularly-symmetric complex Gaussian random vector with mean $\mathbf{\bar{y}}$ and covariance matrix $\mathbf{Z}$ is represented by $\mathbf{y}\sim\mathcal{CN}(\mathbf{\bar{y}},\mathbf{Z})$. Furthermore, $\mathcal{U}(0,1)$ is a uniform distribution between 00 and $1$. Table II lists the symbols used in this paper and their definitions.
II SYSTEM MODEL AND PROBLEM FORMULATION
Consider the IIoT alarm scenario in Fig. 1, where $M$ orthogonal channels are shared by a set $\mathcal{N}$ of wireless IIoT devices to communicate with the BS and the ExC. By utilizing $M$ orthogonal channels, the devices can transmit information concurrently over multiple channels. This significantly improves the likelihood of successful reception of the information at the receiver, in contrast to the scenario in which the devices have access to only a single channel. We assume that ${|\mathcal{N}|\gg M}$ and the devices are time synchronized [6]. At a certain time, an alarm is generated at a random location, called the epicenter. The devices that detect the alarm are called active devices, while the other devices are called inactive devices. Triggered by an alarm event, a random batch of devices ${\mathcal{N}^{\prime}\subseteq\mathcal{N}}$ becomes active. Let ${f(d_{\upsilon})}$ denote the activation probability function of device ${\upsilon\in\mathcal{N}}$. It is a decreasing function of the distance $d_{\upsilon}$ between the device $\upsilon$ and the alarm epicenter [4, 27].
The sole purpose of the active devices is to convey the alarm information to the ExC, which is not necessarily collocated with the BS, as illustrated in Fig. 1. Thus, they must successfully transmit the alarm information on at least one channel, irrespective of which active device makes it possible. For transmitting the alarm information, the active device ${\upsilon\in\mathcal{N}^{\prime}}$ chooses a transmission pattern denoted by ${\boldsymbol{a}_{\upsilon}=[a_{\upsilon,1},a_{\upsilon,2},\cdots,a_{\upsilon,M}]^{T}}$, where
$a_{\upsilon,m}=\left\{\begin{array}[]{ll}1,&\textrm{if device}\ \upsilon\ \textrm{decides to transmit over}\\ &\textrm{channel}\ m,\\ 0,&\textrm{otherwise}.\end{array}\right.$ (4)
We represent the transmission patterns chosen by the active devices as a matrix ${\mathbf{A}\in\{0,1\}^{M\times|\mathcal{N}^{\prime}|}}$, where each column of $\mathbf{A}$ matches with one of the vectors of the set ${\{\boldsymbol{a}_{\upsilon}|\upsilon\in\mathcal{N}^{\prime}\}}$. The successful/failed alarm reception at the ExC is indicated by
$\xi(\mathcal{N}^{\prime},\mathbf{A})=\left\{\begin{array}[]{ll}\!\!1,&\!\!\!\!\textrm{if}\ \exists m\in\{1,...,M\}\!\!:\!\!\displaystyle\sum_{\upsilon\in\mathcal{N}^{\prime}}a_{\upsilon,m}=1,\\ \!\!0,&\!\!\!\!\textrm{otherwise},\end{array}\right.$ (7)
[…]
Figure 2: Illustration of the proposed protocol exemplified for the case of $M\!=$ 2 orthogonal channels. First, each active device transmits a pilot signal to the BS. Then, the BS broadcasts the received pilot signal to the active devices (context broadcast). Finally, the active devices transmit the alarm message after choosing their respective transmission patterns. In the structure of the alarm message, the alarm flag indicates the occurrence of an event, the header carries the metadata, and the payload contains relevant information for the ExC.
With $M$ channels, an active device can choose from $2^{M}$ transmission patterns. We compile all possible transmission patterns into a matrix ${\mathbf{\mathring{A}}=[\boldsymbol{\mathring{a}}_{1},\boldsymbol{\mathring{a}}_{2},\cdots,\boldsymbol{\mathring{a}}_{2^{M}}]\in\{0,1\}^{M\times 2^{M}}}$, where each column represents a unique transmission pattern and $\boldsymbol{\mathring{a}}_{i}\neq\boldsymbol{\mathring{a}}_{j},\forall i\neq j$. For instance, consider the case shown in Fig. 2 where $M=2$, $|\mathcal{N}^{\prime}|=5$, devices $1$, $2$, $4$ select channel $2$, device $5$ selects both channels, and device $3$ is remains silent. Then,
$\mathbf{\mathring{A}}=\left[\begin{matrix}0&0&1&1\\ 0&1&0&1\end{matrix}\right]\textrm{and}\ \mathbf{A}=\left[\begin{matrix}0&0&0&0&1\\ 1&1&0&1&1\end{matrix}\right].$
Therefore, $\xi(\mathcal{N}^{\prime},\mathbf{A})=1$ since $\sum_{\upsilon\in\mathcal{N}^{\prime}}a_{\upsilon,1}=1$.
Let us define the probability of the active device $\upsilon$ choosing a transmission pattern $\boldsymbol{\mathring{a}}_{i}$ as $\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})$. Let the matrix ${\mathbf{\Psi}\in[0,1]^{|\mathcal{N}|\times 2^{M}}}$ store the elements of the following set ${\{\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})|\forall i=1,\cdots,2^{M}}\}$ in its $\upsilon^{th}$ row. Additionally, ${\sum_{i=1}^{2^{M}}\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})=1\textrm{,}\ \forall\upsilon\in\mathcal{N}^{\prime}}$. Now, the expected probability of a successful transmission can be written as [6]
$\!\!\!\Lambda(\mathbf{\Psi})=\!\!\!\!\!\!\sum_{\mathcal{N}^{\prime}\in\mathcal{P}(\mathcal{N})}\prod_{\upsilon\in\mathcal{N}^{\prime}}\!\!f(d_{\upsilon})\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\mathbf{A}\in\{0,1\}^{M\times|\mathcal{N}^{\prime}|}}\!\!\!\!\!\!\!\!\!\!\!\Bigl{(}\xi(\mathcal{N}^{\prime},\mathbf{A})\!\prod_{\upsilon\in\mathcal{N}^{\prime}}\!\!\psi_{\upsilon}(\boldsymbol{a}_{\upsilon})\!\Bigl{)}.$ (8)
The optimization problem for the alarm transmission task can be expressed as
$\begin{split}\underset{\mathbf{\Psi}}{\textrm{maximize}}\ &\Lambda\\ \textrm{subject to}\ &\sum_{i=1}^{2^{M}}\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})=1\textrm{,}\ \forall\upsilon\in\mathcal{N}^{\prime},\\ &\mathbf{\Psi}\in[0,1]^{|\mathcal{N}|\times 2^{M}}.\end{split}$ (9)
[…]
III NEURAL NETWORK-BASED BANDIT
NNBB is a DRL-based RA algorithm that is implemented in each IIoT device. The NNBB components are: (i) an agent (device) that receives a context, (ii) actions (transmission patterns), (iii) an action value ($\psi$) corresponding to each action, and (iv) a reward (feedback signal from the ExC) for updating the action values. The proposed protocol for the alarm transmission is illustrated in Fig. 2. The aforementioned NNBB components and the algorithm implementation are explained in detail below.
Figure 3: DNN Architecture. The input to the DNN is the received context, while its outputs are 2M parameterized action values, one for each possible action.
III-A CONTEXT
An active agent (device) $\upsilon$ observes the context $\boldsymbol{s}_{\upsilon}$. For this, each active agent transmits a pilot signal consisting of $M$ pilot symbols, one on each channel. The BS then receives the signal aggregating all the pilot signals from the active devices and broadcasts it to the agents in the next time slot. The signal $\boldsymbol{s}_{\upsilon}$ received by the active agent $\upsilon$ informs it that some devices are active in the surrounding environment and implicitly captures hidden information about their identity. The goal of the NNBB-based RA mechanism is to make distributed access decisions based on such a context and learning experience.
Assume for simplicity that the BS, ExC, and IIoT devices are single-antenna nodes. Let ${\boldsymbol{\varrho}_{\upsilon}=[\varrho_{1,\upsilon},\dots,\varrho_{M,\upsilon}]^{T}\in\mathbb{C}^{M\times 1}}$ represent the pilot sequence for an active agent $\upsilon$. Therefore, the aggregated pilot signal $\boldsymbol{s}$ received by the BS can be written as
$\boldsymbol{s}=\sum_{\upsilon\in\mathcal{N}^{\prime}}\sqrt{\rho}\operatorname{diag}(\mathbf{c}_{\upsilon})\boldsymbol{\varrho}_{\upsilon}+\boldsymbol{\varphi},$ (10)
where $\rho$ represents the average transmit signal-to-noise ratio, ${\boldsymbol{\varphi}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})\in\mathbb{C}^{M\times 1}}$ is the Additive White Gaussian Noise (AWGN) with normalized power, and ${\mathbf{c}_{\upsilon}\in\mathbb{C}^{M\times 1}}$ captures the $M$ channel coefficients between the device $\upsilon$ and the BS. After receiving $\boldsymbol{s}$, the BS broadcasts $\boldsymbol{s}$ to the devices. Therefore, the context $\boldsymbol{s}_{\upsilon}$ received by the active device $\upsilon$ can be written as
$\boldsymbol{s}_{\upsilon}=\sqrt{\rho}\operatorname{diag}(\mathbf{c}_{\upsilon})\boldsymbol{s}+\boldsymbol{\hat{\varphi}},$ (11)
where ${\boldsymbol{\hat{\varphi}}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})\in\mathbb{C}^{M\times 1}}$ is the AWGN with normalized power. Here, we assume that $\rho$ is the same as the one used for pilot transmission to simplify the analysis without any loss of generality.
III-B ACTION
After receiving $\boldsymbol{s}_{\upsilon}$, the active agent $\upsilon$ selects a transmission pattern $\boldsymbol{a}_{\upsilon}$, which we shall refer to as its action, and performs a single-hop alarm transmission to the ExC. Recall that with $M$ channels, an active device can choose from $2^{M}$ actions, which are collected as columns of the matrix $\boldsymbol{\mathring{A}}$. The alarm message consists of a header, alarm flag, and payload as shown in Fig. 2. The alarm flag indicates the occurrence of an event, the header carries the metadata, and the payload contains relevant information about the alarm for the ExC.
III-C REWARD
The active devices’ cumulative target is to successfully transmit alarm information on at least one channel. Therefore, the alarm message transmission is considered successful when the condition ${\xi(\mathcal{N}^{\prime},\mathbf{A})=1}$ is met. When an alarm message transmission succeeds, all active agents receive a shared Acknowledgement (ACK) from the ExC, in which case the active agent $\upsilon$ gets a reward ${r(\boldsymbol{a}_{\upsilon})=1}$. Otherwise, ${r(\boldsymbol{a}_{\upsilon})=0}$.
III-D ACTION VALUE AND SELECTION
An agent aims to maximize its total reward obtained over time. To achieve this, it is attractive to choose those actions that have yielded a reward in the past. This is known as exploitation. However, to discover such reward-accruing actions, an agent must try new actions, which is known as exploration. In our work, an agent balances exploration and exploitation using the $\varepsilon$-greedy method. Such a method relies on the action value ${q(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i}),\ \forall i=1,\cdots,2^{M}}$ that represents a prediction of the expected reward for agent $\upsilon$ when it takes an action $\boldsymbol{\mathring{a}}_{i}$ with context $\boldsymbol{s}_{\upsilon}$. In the $\varepsilon$-greedy method, an agent $\upsilon$ primarily selects a greedy action (exploitation) ${\boldsymbol{a}_{\upsilon}=\operatorname{argmax}_{\boldsymbol{\mathring{a}}_{i}\in\{0,1\}^{M}}q(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i})}$, with occasional exploration by randomly selecting an action $\boldsymbol{a}_{\upsilon}$ from all possible actions with a probability $\varepsilon$. In this way, every action is eventually selected many times. It is necessary to decrease $\varepsilon$ to gradually emphasize exploitation over exploration. We gradually decrease $\varepsilon$ from $1$ to $0.1$ with a step size of $0.005$ after every alarm event.
Due to channel fading and noise, the number of possible contexts is infinite, leading to an agent encountering a new context every time it becomes active. Thus, maintaining separate action values for each context is not a reasonable approach. Rather, the agent should maintain the action values as a parameterized function and tune its parameters to better match $q$ with the reward obtained after observing the feedback from the ExC. The parameterized function, for which a DNN is employed, allows an estimation of the action value without maintaining a table of action values.
The parameterized action-value function for the agent $\upsilon$ is expressed as ${\hat{q}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i},\mathbf{w})}$, where $\mathbf{w}$ is a vector of connection weights present in the DNN layers. The input to the DNN is the received context, as illustrated in Fig. 3. The DNN comprises two fully connected hidden layers and a dense output layer. Notice that two hidden layers are capable of approximating any smooth mapping with arbitrary accuracy [28]. Also, to grapple with the exploding gradients, we use gradient norm clipping [29]. The outputs from the DNN are $2^{M}$ parameterized action values $\hat{q}$. In addition, the DNN employs (a) ReLU non-linearity, which is a computationally efficient thresholding operation, compared to sigmoid and tanh, that helps alleviate the problem of vanishing gradients, and (b) RMSProp optimizer that maintains a moving average of the squared gradient values for each weight to adjust the update size for each weight. Moreover, the learning rate of the DNN is a hyper-parameter in RMSProp; we decrease it with a decay rate of $0.015$ per alarm event.
After calculating the reward ${r(\boldsymbol{a}_{\upsilon})}$, the active agent $\upsilon$ stores the tuple ${\{\boldsymbol{s}_{\upsilon},\boldsymbol{a}_{\upsilon},r(\boldsymbol{a}_{\upsilon})\}}$ in its memory buffer $E$. If $E$ is full, it removes an already stored tuple from $E$ in a First-In-First-Out (FIFO) fashion. Indeed, there is no need to store the entire dataset in the agent’s memory.
The active agent $\upsilon$ trains its DNN in an online manner. Training is performed by sampling a mini-batch of size $B$ from $E$ and providing this mini-batch as input to the DNN to update $\mathbf{w}$ by minimizing the loss function
$\Omega=\frac{1}{B}\sum_{j=1}^{B}[r_{j}(\boldsymbol{a}_{\upsilon})-\hat{q}_{j}(\boldsymbol{s}_{\upsilon},\boldsymbol{a}_{\upsilon},\mathbf{w})]^{2}$ (12)
using RMSProp. Herein, $r_{j}$ and $\hat{q}_{j}$ represent the reward and parameterized action value, respectively, for the $j^{th}$ sample in the mini-batch. Meanwhile, the gradient vector $\nabla_{\mathbf{w}}\Omega$ is clipped in the following manner:
$\boldsymbol{\chi}\ =\ \frac{\beta_{0}\ \nabla_{\mathbf{w}}\Omega}{\max({\|\nabla_{\mathbf{w}}\Omega\|}_{2},\beta_{0})},$ (13)
where $\beta_{0}$ represents the threshold value for ${\|\nabla_{\mathbf{w}}\Omega\|}_{2}$ and the vector $\boldsymbol{\chi}$ stores the clipped gradients.
III-E ALGORITHM IMPLEMENTATION
The NNBB is implemented in each IIoT device detecting an alarm event. Such active devices execute the following operations:
  1. 1.
    Transmit their respective pilot signals to the BS, which then broadcasts the received aggregated signal back to them. Consequently, each active agent acquires its context.
  2. 2.
    Feed their context to their DNN, which yields $2^{M}$ action values.
  3. 3.
    Select an action using the $\varepsilon$-greedy method and transmit the alarm message accordingly.
  4. 4.
    Monitor the feedback from the ExC and determine its reward.
  5. 5.
    Store the received context, the selected action, and the acquired reward in their memory.
  6. 6.
    Train their DNN.
Algorithm 1 provides a comprehensive description of NNBB.
Algorithm 1 NNBB at active agent $\upsilon$
0.  Context $\boldsymbol{s}_{\upsilon}$, exploration rate $\varepsilon$, $\mathbf{w}$
1.  Evaluate ${\hat{q}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i},\mathbf{w})\ \forall i=1,\cdots,2^{M}}$ using the DNN
2.  Draw $\Theta\in\mathcal{U}(0,1)$
3.  if $\Theta>\varepsilon$ then
4.     $\boldsymbol{a}_{\upsilon}=\operatorname{argmax}_{\boldsymbol{\mathring{a}}_{i}\in\{0,1\}^{M}}\hat{q}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i},\mathbf{w})$
5.  else
6.     Select $\boldsymbol{a}_{\upsilon}$ randomly
7.  end if
8.  if $E$ is full then
9.     Remove a tuple from $E$ in a FIFO fashion
10.  end if
11.  Observe the feedback from the ExC. If an ACK is received then $r(\boldsymbol{a}_{\upsilon})=1$, otherwise $r(\boldsymbol{a}_{\upsilon})=0$
12.  Store the tuple ${\{\boldsymbol{s}_{\upsilon},\boldsymbol{a}_{\upsilon},r(\boldsymbol{a}_{\upsilon})\}}$ in $E$
13.  Sample a mini-batch of size $B$ from $E$
14.  Provide this mini-batch as input to the DNN for updating $\mathbf{w}$ by minimizing $\Omega$ using RMSProp
15.  $\varepsilon\leftarrow\max(0.1,\ \varepsilon-0.005)$
15.  Transmission pattern $\boldsymbol{a}_{\upsilon}$, $\varepsilon$, $\mathbf{w}$
III-F MAC IN NNBB
The MAC of an IIoT device in NNBB is described by three states: Normal State (NS), Emergency State (ES), and Quiet State (QS). In NS, the IIoT device transmits data from the sensed process. In ES, an active device transmits the alarm message. In QS, an inactive device halts its process data transmission. The latter can be resumed in a future time slot, provided the device is no longer in QS.
Fig. 4 illustrates the MAC procedure followed by active and inactive devices after the occurrence of an alarm event in a time slot. During phase-1, the active devices transition to ES while the inactive devices stay in NS. Next, in phase-2, the active devices transmit pilot signals, and the inactive devices scan available channels to determine if they are idle or busy. If all channels are idle, the inactive devices proceed with the transmission of their process data in the subsequent phase. If all channels are busy, the inactive devices enter QS. Following phase-2, the active devices go through phase-3a, phase-3b, and phase-4, where they receive context, transmit alarm messages, and receive ACK/no-ACK, respectively. Here, no-ACK indicates an unsuccessful alarm transmission. Meanwhile, during phase-3, and phase-4, the inactive devices stay in QS. At last, in phase-5, both active and inactive devices transition to NS.
Note that our target is to improve the success rate of the first transmission attempt of the alarm message, rather than focusing on its retransmissions. Thus, we skipped the alarm retransmission procedure.
Figure 4: Illustration of the MAC procedure followed by an active and inactive device after the occurrence of an alarm event.
TABLE III: Parameters Used in Simulations
Parameters Value
Density of the circular region $0.2$ devices///m2
Initial learning rate for DNN and MAB $1.0$
Total number of algorithm runs $100$
Mini-batch size $(B)$ $2^{M}\times 30$
Memory buffer size $(E)$ [30] $2^{M}\times 100$
Number of hidden layers $(H)$ $2$
Size of each hidden layer $(h)$ $1$
Global norm (of gradient vector) threshold $(\beta_{0})$ [30] $5.0$
Path loss exponent $(\gamma)$ $3.8$
Mean-scaling multiplier $(\lambda)$ $3$
IV NUMERICAL RESULTS
We now assess the proposed NNBB RA scheme through simulations. We consider a circular region around the BS, where the ExC and the devices are uniformly distributed. The circular region has a fixed device density of $0.2$ devices///m2, and we vary the radius of the area to accommodate the specified number of devices. Similar to [4], we use
$f(d_{\upsilon})=e^{-d_{\upsilon}/\lambda},$ (14)
as the activation probability function, where $\lambda$ is a mean-scaling multiplier. The channels undergo Rayleigh fading and ${\mathbf{c}_{\upsilon}\sim\mathcal{CN}(\mathbf{0},\frac{1}{r_{\upsilon}^{\gamma}}\mathbf{I}_{M})}$, where ${r_{\upsilon}^{\gamma}}$ represents the path loss, $\gamma$ is the path loss exponent, and $r_{\upsilon}$ is the distance between the device $\upsilon$ and the BS. The simulation parameters and their corresponding values are listed in Table III. We measure the performance of a RA scheme using the success rate metric, which is defined as the probability that an alarm event is successfully reported to the ExC. We measure the success rate of an algorithm after it has converged. To illustrate the variability in the success rate of the first transmission attempt in our simulation results, we generate an alarm at each time slot and discard the respective alarm message from the active devices after its first transmission attempt.
IV-A BENCHMARK SCHEMES
We use MAB-based RA [31] and RS as benchmarks. Akin to NNBB, MAB and RS are implemented at each IIoT device. MAB uses the $\varepsilon$-greedy method to select an action when an alarm event is detected. The action values, $Q(\boldsymbol{\mathring{a}}_{i})\ \forall i=1,\cdots,2^{M}$, in the MAB-based RA scheme are updated as
$Q(\boldsymbol{\mathring{a}}_{i})\leftarrow(1-\tau)Q(\boldsymbol{\mathring{a}}_{i})+r(\boldsymbol{\mathring{a}}_{i})\tau,$ (15)
where $\tau$ is the learning rate with an initial value of $1$ and it decreases with time. In contrast with [6], we simultaneously train the MABs of all the active devices at each alarm event, i.e., the action values of every active device are updated. On the other hand, in the RS scheme, actions are selected randomly at each alarm event.
Figure 5: Comparison of the success rates of various RA algorithms for different numbers of available channels in the network.
IV-B IMPACT OF THE NUMBER OF DEVICES AND CHANNELS
Fig. 5 shows the impact of the number of devices on the success rate achieved by NNBB and the benchmark schemes. As the number of devices increases, the success rate deteriorates. However, NNBB increasingly outperforms MAB and RS. Specifically, for a network of $20$ devices with $5$ channels, the success rate of NNBB is $9\%$ higher than that of MAB. This gain increases to $20\%$ when there are $60$ devices. Although MAB outperforms NNBB when only $2$ channels are shared by at most $20$ devices, the superiority is marginal. The superiority disappears and gets reversed as the number of devices increases. All in all, as the number of devices increases, the performance gain of NNBB over MAB and RS increases.
Fig. 6 plots the success rate as a function of the number of channels. We observe that NNBB outperforms MAB and RS regardless of the number of available channels for heavily loaded networks. Notably, increasing the number of available channels from $2$ to $6$ results in a $30\%$ increase in the success rate for NNBB in a network of $60$ devices, whereas MAB only experiences a $13\%$ improvement. A closer examination of the plots in Fig. 6 also reveals that NNBB is less sensitive to an increase in the number of devices than MAB and RS.
Figure 6: Comparison of the success rates of various RA algorithms for different numbers of devices in the network.
Figure 7: Comparison of the success rates of various RA algorithms as a function of the scaling multipliers $\lambda$, when the network consists of 5 channels.
IV-C IMPACT OF THE ACTIVATION PROBABILITY
Fig. 7 illustrates the effect of $\lambda$, which directly impacts the activation probability, on the performance of the RA schemes. As $\lambda$ increases, the activation probability increases. The figure indicates that NNBB and the benchmark schemes exhibit similar success rates when the activation probability is small, i.e., when $\lambda=1$. However, as the activation probability increases, the success rate of NNBB decays at a slower rate compared to MAB and RS. Even for $\lambda=3$, when the success rates of the benchmark schemes fall well below $0.8$, the success rate of NNBB remains significantly higher. Notably, for NNBB, a significant difference in the success rates between the two considered networks is visible only when $\lambda=4$, while for MAB and RS the difference becomes apparent already when $\lambda=3$.
IV-D IMPACT OF THE NUMBER AND SIZE OF THE HIDDEN LAYERS
Figure 8: Comparison of the success rates of NNBB for various combinations of hidden layers, when the network consists of 40 devices.
Fig. 8 illustrates the performance of NNBB as a function of the number of hidden layers and their size. It is evident from the figure that NNBB consistently outperforms the benchmark approaches. Interestingly, the hidden layer combination with the least number of neurons, i.e., ${(H,h)=(2,1)}$, yields the highest success rate among the considered configurations. Additionally, this configuration outperforms the other two configurations throughout the training procedure, as indicated in Fig. 9. It is often advantageous to reduce the hidden layer size as much as possible to maintain the NN’s generalization capability. The excess neurons can cause the network to act like a memory bank, leading to suboptimal performance when presented with inputs other than the training samples [32]. This justifies the suboptimal performance of the configurations ${(H,h)=(1,10)}$ and ${(H,h)=(2,15)}$. Reducing the hidden layer size also reduces the computational complexity of NNBB. Furthermore, Fig. 9 demonstrates that for a network of $40$ devices, NNBB should train over $15200$ and $31000$ alarm events to converge when $5$ and $6$ channels, respectively, are available.
Figure 9: Comparison of the performance of NNBB (during training) for various combinations of hidden layers, when the network consists of 40 devices and (a) 6 channels (left) or (b) 5 channels (right).
V CONCLUSION
We proposed NNBB, which is a distributed DRL-based RA scheme that allows IIoT devices to develop implicit coordination to successfully convey an alarm message to an external controller. Specifically, upon the detection of an alarm event, every active device starts a procedure to acquire a useful context to feed a local DNN. Then, with the help of the DNN and the $\varepsilon$-greedy method, a device selects the transmit channel(s) for the alarm message, including also the possibility of no transmission. A reward or penalty is granted based on the success or failure of the transmission, which eventually is used for training the DNN. Notably, the DNN uses two hidden layers with just one neuron each, significantly decreasing NNBB’s computational complexity. Simulations show that NNBB experiences a relatively lower drop in its success rate compared to the benchmark schemes as the number of devices in a network increases. Further, the success rate gained by increasing the available channels is relatively higher.
For future work, it might be interesting substituting the adopted $\varepsilon$-greedy method, which is less effective in large action spaces and sensitive to the initial value of $\varepsilon$, by 1) Thompson sampling-based exploration, and 2) upper confidence bound-based exploration. Meanwhile, NNBB with the help of context develops implicit coordination among devices. Therefore, exploiting NNBB in connected robotics and autonomous systems are interesting future research directions.